Session: Feb 2011 (session 1)
Wavelength: 18 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (Franco Fiocchi at Sun Feb 27 17:12:38 2011) Success:
Light rain

Onsala (Michael Lindqvist at Sun Feb 27 18:07:43 2011) Success:
No known problems.

Westerbork (Geert Kuper at Mon Feb 28 07:37:13 2011) Success:
Only one telescope was used for this project, as requested by the PI. (Only telescope 7 in the adding section.)

Torun (K. Borkowski at Mon Feb 28 08:49:37 2011) Success:
No known problems.

Arecibo (Tapasi Ghosh at Tue Mar 01 13:31:30 2011) Success:

Effelsberg (Uwe Bach at Tue Mar 01 14:33:18 2011) Minor failures:
Lost 3 scans at 13 UT because of antenna tracking problems.

Jodrell_Bank (Alastair Gunn at Tue Mar 29 11:27:12 2011) Minor failures: ANT
Several minutes lost at 12:03 and 14:49 due to control problem. No other known problems.